# Understanding Multimodal LLMs Under Distribution Shifts: An Information-Theoretic Approach

### v1

@@ 4.3 Characterizing MLLM Performance Gap via Effective Mutual Information Difference
${\small\begin{split}\text{EMID}(P_{\mathbf{X}Y},Q_{\mathbf{X}Y};\theta):=\text{EMI}(P_{\mathbf{X}Y};\theta)-\text{EMI}(Q_{\mathbf{X}Y};\theta).\end{split}}$ (9)
To elucidate the key insight and provide a clear foundation, we begin by analyzing a simple scenario where the conditional variables remain consistent across both ID and OOD distributions. In this case, we can derive an upper bound for EMID, as stated in Theorem 4.5. This bound enables us to quantify the maximum performance gap of MLLM over two distributions by measuring the severity of the marginal distribution shift over visual and language modalities.
@@
and denote $P_{Y_{\theta}}=\mathbb{E}_{P_{\mathbf{X}}}[P_{{\theta}}(\cdot|\mathbf{x})]$ and $Q_{Y_{\theta}}=\mathbb{E}_{Q_{\mathbf{X}}}[P_{{\theta}}(\cdot|\mathbf{x})]$, then $\text{EMID}(P_{\mathbf{X}Y},Q_{\mathbf{X}Y};\theta)$ is upper bounded by
@@
Figure 3: Scatter plot with regression line between empirical estimates of EMID and its upper bound. Over the 34 synthetic and 27 natural distribution shift scenarios, we evaluate four MLLMs and get 136 cases and 108 cases of synthetic shifts and natural shifts, respectively, for visualizing EMID and upper bound estimates. The two panels on the left show results for all four models, whereas the right ones distinguish them per model with fitter linear regression coefficients (Slope).
@@
Theorem 4.5 naturally captures special cases such as visual-only or text-only input shifts. For a visual-only input shift, where $D_{\rm JS}(P_{X_{t}}\|Q_{X_{t}})=0$, the EMID upper bound primarily depends on the divergence between the visual input distributions. Similarly, for a text-only input shift, the bound reflects the divergence in the textual input distributions. The two cases not only underscore the flexibility of Theorem 4.5 in isolating and quantifying the impact of modality-specific distribution shifts on model performance but also highlight the importance of visual and text input shifts on it. In Appendix D, we further provide a looser yet better interpretable version of this upper bound (Corollary D.12) by replacing the $\Delta$ into the discrepancy terms between model output and ground truth conditional distributions.
@@
Moving beyond the simplified scenario, we now consider the general scenario where no assumptions are made about the consistency of conditional distributions across ID and OOD settings. This more realistic scenario accommodates shifts not only in the marginal distributions of visual and textual inputs but also in their conditional dependencies and the relationships between inputs and outputs. By relaxing these constraints, we aim to capture the full complexity of distributional shifts encountered in practice and analyze how such shifts collectively influence the performance gap of MLLMs. The formal upper bound is provided in Theorem 4.6.
@@
and denote $P_{Y_{\theta}}=\mathbb{E}_{P_{\mathbf{X}}}[P_{{\theta}}(\cdot|\mathbf{x})]$ and $Q_{Y_{\theta}}=\mathbb{E}_{Q_{\mathbf{X}}}[P_{{\theta}}(\cdot|\mathbf{x})]$, then $\text{EMID}(P_{\mathbf{X}Y},Q_{\mathbf{X}Y};\theta)$ is upper bounded by
@@
where $\widehat{H}=\max_{\mathbf{x}\in\mathcal{X}}[H(Q_{Y|\mathbf{X}=\mathbf{x}})+H(P_{\theta}(\cdot|\mathbf{x}))]$, $\Delta=\delta_{P}+\delta_{Q}$, and
@@
Although Theorem 4.6 holds for broader cases, Theorem 4.5 is much simpler to analyze. Thus, we focus on the validation of Theorem 4.5 in the following section. If we have some knowledge of the data-generating process of $P_{\mathbf{X}Y}$ and $Q_{\mathbf{X}Y}$, we can choose the one that is suitable for given distributions. In summary, both Theorem 4.5 and 4.6 provide an analytic tool to characterize the performance gap of MLLM, representing the first formal framework for evaluating MLLM under distribution shifts.
@@ 5 Empirical Validation on Real Benchmark
As done in the pilot experiments (Figure 1 and 6), we used LLaVA v1.5 (Liu et al., 2024a) and LLaVA NeXT (Liu et al., 2024b) in 7B and 13B sizes and evaluated them on the LLaVA-Bench COCO and LLaVA-Bench Wild (Liu et al., 2023) datasets for assessing open-ended generation. To comprehensively examine diverse types of shifts, we further simulate synthetic distribution shifts as well as natural distribution shifts. For synthetic shifts, we consider 7 visual scenarios (1 ID case $+$ 2 synthetic perturbation types at 3 severity levels), and 5 text scenarios (1 ID case $+$ 2 synthetic perturbation types at 2 severity levels), resulting in $7\times 5=35$ synthetic scenarios, where 1 scenario is ID and the other 34 are OOD cases. For natural shifts, we use 4 visual scenarios (1 ID + 3 OOD difficulty levels) and 7 text scenarios (1 ID-English $+$ 6 languages), yielding $4\times 7=28$ natural scenarios. This comprehensive design covers a total of 34 synthetic and 27 natural shifts. A summary of the OOD construction strategies is in Table 1.
@@
(in-the-wild Images)
LLaVA-Bench in-the-wild
@@
For the empirical realization of our theoretical statements, we adopt a popular neural estimator for MI, CLUB (Cheng et al., 2020) to compute empirical EMID, and a JS divergence estimator, RJSD (Hoyos-Osorio & Sanchez-Giraldo, 2023), to compute empirical EMID and its upper-bound (see Appendix B for details). Experiments with 23 alternative implementations derived consistent conclusions (Please refer to Table 4 and 5).
Correlation between win rate and EMI.
We first conduct the Spearman correlation analysis and Kendall’s tau analysis between the win rate and our empirical estimates of EMI. In Table 2, we can see that EMI estimates exhibit a strong correlation with win rate, both in terms of absolute coefficient and $p$-value, across all models. This empirical evidence validates the theoretical connection between EMI and win rate discussed in Theorem 4.4. Therefore, our EMI can be used as a reliable and cost-efficient alternative to win rate for MLLM evaluation with theoretical guarantees.
Table 2: Spearman rank correlation and Kendall’s tau between win rate and EMI. We conduct correlation analysis between the win rate (Eq. (2)) and EMI (Eq. (7)) on 34 synthetic and 27 natural distribution shifts across four MLLMs.
@@
Table 3: Pearson correlation analysis between EMID and its upper bound. We provide Pearson $r$ and $p$-value between the empirical estimates of EMID (Eq. (9)) and its upper bound (Eq. (10)) on 34 synthetic and 27 natural shifts scenarios per model.
@@
LLaVA v1.5 7B 0.742 $<$0.001 0.594 0.001
LLaVA v1.5 13B 0.807 $<$0.001 0.550 0.003
@@
We now validate our main Theorem. Figure 3 (left two) shows the scatter plots comparing EMID with its upper bound across four models, each evaluated over 34 and 27 synthetic and natural distribution shifts. We see a clear trend between EMID and its upper bound in the synthetic shift where we could directly control the severity of shifts. While the natural shift setup is noisier, a similar overall trend is observed. Meanwhile, our bounds depend on the distributional discrepancy between the model’s response $\hat{y}$ and the ground truth response $y$. Thus, they naturally induce different bounds for each MLLM. The right panel of Figure 3 presents model-wise plots with linear regression coefficients, where we observe that each model has a different degree of performance sensitivity against shifts. Pearson correlation analysis results in Table 3 further confirm statistically significant correlations between EMID and its upper bound, supporting the validity of the theorem.
@@
It is common that we can not access the ground truth response $Y$ from our evaluation dataset in advance. Then, one may want to use the EMID upper bound as an estimator of the maximum risk of MLLM given two datasets, i.e., $\max\text{EMID}(P_{\mathbf{X}Y},Q_{\mathbf{X}Y};\theta)$, by neglecting the output-related term $\Delta$. In Figure 4, we investigate whether the summation of two JS divergence terms can still be predictive for EMID. Although the trends become loose compared to the full bound due to the non-optimality of MLLM parameters, the partial upper bound still has moderate correlations (denoted by Pearson $r$) with EMID.
@@ 6 Related Works
Recent findings imply that fine-tuning on a relatively small amount of ID datasets hurts the OOD generalization capability of foundation models (Kumar et al., 2022; Wortsman et al., 2022). Although lots of follow-up studies (Goyal et al., 2023; Tian et al., 2023; Oh et al., 2025b) including theory-inspired methods (Kumar et al., 2022; Ju et al., 2022; Oh et al., 2025a) have been proposed, almost all of them focused on a discriminative model such as CLIP (Radford et al., 2021) for the image classification task. Given the rapidly growing popularity of MLLMs, it is necessary to investigate the reliability of MLLMs under distribution shifts with a tangible formulation. We lay a cornerstone for this.
@@
As well as learning objectives (Alemi et al., 2016; Chen et al., 2016; Tschannen et al., 2020; Kong et al., 2020; Wang et al., 2021), information-theoretic view has been steadily adopted to establish evaluation criteria for language model probing (Hewitt et al., 2021), prompt engineering (Sorensen et al., 2022), and rationale evaluation (Chen et al., 2023), but relatively unexplored for MLLM yet. We also note some works adopting information-theoretic approaches to analyze models under distribution shifts (Federici et al., 2021; Shui et al., 2022). Although they focused on classification tasks with discriminative models, we established new theorems for MLLM analysis based on our new metric, EMI.

### v2

@@ 4.3 Characterizing MLLM Performance Gap via Effective Mutual Information Difference
${\small\begin{split}\text{EMID}(P_{\mathbf{X}Y},Q_{\mathbf{X}Y};P_{\theta}):=\text{EMI}(P_{\mathbf{X}Y};P_{\theta})-\text{EMI}(Q_{\mathbf{X}Y};P_{\theta}).\end{split}}$ (9)
To elucidate the key insight and provide a clear foundation, we begin by analyzing a simple scenario where the conditional variables remain consistent across both ID and OOD distributions. In this case, we can derive an upper bound for EMID, as stated in Theorem 4.5. This bound enables us to characterize the maximum performance gap of MLLM over two distributions by measuring the severity of the marginal distribution shift over visual and language modalities.
@@
$D_{\rm JS}(P_{Y_{{\theta}}}\|P_{Y})\leq\delta_{P},~{}~{}~{}D_{\rm JS}(Q_{Y_{{\theta}}}\|Q_{Y})\leq\delta_{Q},\;\;\Delta=\delta_{P}+\delta_{Q}$
and denote $P_{Y_{\theta}}=\mathbb{E}_{P_{\mathbf{X}}}[P_{{\theta}}(\cdot|\mathbf{x})]$ and $Q_{Y_{\theta}}=\mathbb{E}_{Q_{\mathbf{X}}}[P_{{\theta}}(\cdot|\mathbf{x})]$, then $\text{EMID}(P_{\mathbf{X}Y},Q_{\mathbf{X}Y};P_{\theta})$ is upper bounded by
@@
where $\widehat{H}=\max_{\mathbf{x}\in\mathcal{X}}[H(Q_{Y|\mathbf{X}=\mathbf{x}})+H(P_{\theta}(\cdot|\mathbf{x}))]$.
Figure 3: Scatter plot with regression line between empirical estimates of EMID and its upper bound. Over the 34 synthetic and 27 natural distribution shift scenarios, we evaluate four MLLMs and get 136 cases and 108 cases of synthetic shifts and natural shifts, respectively, for visualizing EMID and its scale-adjusted upper bound estimates (See Appendix B for details). The two panels on the left show results for all four models, whereas the right ones distinguish them per model with fitted linear regression coefficients (Slope).
@@
Theorem 4.5 naturally captures special cases such as visual-only or text-only input shifts. For a visual-only shift, where $D_{\rm JS}(P_{X_{t}}\|Q_{X_{t}})=0$, the EMID upper bound depends primarily on the divergence between visual input distributions, and output discrepancy terms. Similarly, for a text-only shift, the bound reflects the divergence in textual input distributions, and output discrepancy terms. These cases not only underscore the flexibility of Theorem 4.5 in isolating the impact of modality-specific distribution shifts on model performance but also highlight the importance of visual and text input shifts on it. In Appendix D, we provide a looser yet better interpretable version of this upper bound (Corollary D.12) by replacing $\Delta$ with discrepancy terms between model output and ground truth conditional distributions.
@@
Moving beyond the simplified scenario, we now consider the general scenario in which no assumptions are made about the consistency of conditional distributions across ID and OOD settings. This more realistic scenario accommodates shifts not only in the marginal distributions of visual and textual inputs but also in their conditional dependencies and the relationships between inputs and outputs. By relaxing these constraints, we aim to capture the full complexity of distributional shifts encountered in practice and analyze how such shifts collectively influence the performance gap of MLLMs. The formal upper bound is provided in Theorem 4.6.
@@
$D_{\rm JS}(P_{Y_{{\theta}}}\|P_{Y})\leq\delta_{P},~{}~{}~{}D_{\rm JS}(Q_{Y_{{\theta}}}\|Q_{Y})\leq\delta_{Q},\;\;\Delta=\delta_{P}+\delta_{Q}$
and denote $P_{Y_{\theta}}=\mathbb{E}_{P_{\mathbf{X}}}[P_{{\theta}}(\cdot|\mathbf{x})]$ and $Q_{Y_{\theta}}=\mathbb{E}_{Q_{\mathbf{X}}}[P_{{\theta}}(\cdot|\mathbf{x})]$, then $\text{EMID}(P_{\mathbf{X}Y},Q_{\mathbf{X}Y};P_{\theta})$ is upper bounded by
@@
where $\widehat{H}=\max_{\mathbf{x}\in\mathcal{X}}[H(Q_{Y|\mathbf{X}=\mathbf{x}})+H(P_{\theta}(\cdot|\mathbf{x}))]$ and
@@
Theorem 4.6 holds for broader cases, whereas Theorem 4.5 is much simpler to analyze. Thus, we focus on the validation of Theorem 4.5 in the following section. If we have some knowledge of the data-generating process of $P_{\mathbf{X}Y}$ and $Q_{\mathbf{X}Y}$, we can choose the one that is suitable for given distributions. Although the bounds in both theorems are not such tight, they help us to understand the source of the MLLM performance gap. That is, both Theorem 4.5 and 4.6 provide an analytic tool to characterize the performance gap of MLLM, representing the first formal framework for evaluating MLLM under distribution shifts.
@@ 5 Empirical Validation on Real Benchmark
As done in the pilot experiments (Figure 1 and 6), we mainly used LLaVA v1.5 (Liu et al., 2024a) and LLaVA NeXT (Liu et al., 2024b) in 7B and 13B sizes and evaluated them on the LLaVA-Bench COCO and LLaVA-Bench Wild (Liu et al., 2023) datasets to assess open-ended generation quality. We also considered two advanced MLLMs: Qwen2.5-VL-7B (Bai et al., 2025) and InternVL2.5-7B (Chen et al., 2024), and a domain-specialized open-ended question answering dataset, LLaVA-Med (Li et al., 2024), to explore the broad applicability of our framework. For a comprehensive examination on diverse types of shifts, we further simulate synthetic distribution shifts as well as natural distribution shifts. For synthetic shifts, we consider 7 visual scenarios (1 ID case $+$ 2 synthetic perturbation types at 3 severity levels), and 5 text scenarios (1 ID case $+$ 2 synthetic perturbation types at 2 severity levels), resulting in $7\times 5=35$ synthetic scenarios, where 1 scenario is ID and the other 34 are OOD cases. For natural shifts, we use 4 visual scenarios (1 ID + 3 OOD difficulty levels) and 7 text scenarios (1 ID-English $+$ 6 different languages), yielding $4\times 7=28$ natural scenarios. This comprehensive design covers 34 synthetic and 27 natural shifts spanning 61 shift scenarios in total. A summary of the OOD construction strategies is in Table 1.
@@
(Wild Images)
LLaVA-Bench Wild or LLaVA-Med
@@
For the empirical realization of our theoretical statements, we adopt a popular neural estimator for MI, CLUB (Cheng et al., 2020) to compute empirical EMI and EMID, and a JS divergence estimator, RJSD (Hoyos-Osorio & Sanchez-Giraldo, 2023), to compute empirical estimates of EMID upper bound on top of embeddings of $\mathbf{X}$, $Y$, and $\hat{Y}$, extracted from CLIP-ViT-B/32 (Radford et al., 2021) and XLM-RoBERTa-base (Conneau, 2019) (see Appendix B for details). Experiments with 23 alternative implementations with other estimators derived consistent conclusions (Please refer to Table 7 and 8).
Correlation between RP score and EMI.
We first conduct the Spearman correlation analysis and Kendall’s tau analysis between the RP score and our empirical estimates of EMI. In Table 2, we can see that EMI estimates exhibit a strong correlation with RP score, both in terms of absolute coefficient and $p$-value, across all models. This empirical evidence validates the theoretical connection between EMI and RP score discussed in Theorem 4.4. Therefore, our EMI can be used as a reliable and cost-effective alternative to RP for MLLM evaluation with theoretical grounding.
Table 2: Spearman rank correlation and Kendall’s tau between relative preference (RP) score and EMI. We conduct correlation analysis between the RP score (Eq. (2)) and EMI (Eq. (7)) on 34 synthetic and 27 natural distribution shifts across four MLLMs.
@@
Table 3: Pearson correlation analysis between EMID and its upper bound. We provide Pearson $r$ and $p$-value between the empirical estimates of EMID (Eq. (9)) and its upper bound (Eq. (10)) on 34 synthetic and 27 natural shift scenarios per model.
@@
LLaVA NeXT 7B 0.742 $<$0.001 0.594 0.001
LLaVA NeXT 13B 0.807 $<$0.001 0.550 0.003
@@
We now validate our main Theorem. Figure 3 (left two) shows the scatter plots comparing EMID with its upper bound across four models, each evaluated over 34 and 27 synthetic and natural distribution shifts. We see a clear trend between EMID and its upper bound in the synthetic shift where we could directly control the severity of shifts. While the result of the natural shift case is noisier, the overall trend is similar. Meanwhile, our bounds depend on the distributional discrepancy between the model’s response $\hat{y}$ and the ground truth response $y$. Thus, they naturally induce different bounds per MLLM. The right panel of Figure 3 presents model-wise plots with linear regression coefficients, where we observe that each model has a different degree of sensitivity against shifts. Pearson correlation analysis results in Table 3 further confirm statistically significant correlations between EMID and its upper bound, supporting the validity of our theorems.
@@
It is common that we can not access the ground truth response $Y$ from our training and evaluation datasets. Then, one may want to use a part of EMID upper bound (e.g., JSD terms on visual and textual input only) as an estimator of EMID, by neglecting the output-related term $\Delta$. In Figure 4, we investigate whether the summation of two JS divergence terms can still be predictive for EMID. Although the trends become loose compared to the full bound due to the non-optimality of MLLM parameters, the partial upper bound still has moderately high correlations (denoted by Pearson $r$) with EMID.
@@
Validation with advanced MLLMs.
So far, our evaluation has focused on LLaVA series MLLMs. Now we validate EMI and EMID upper bound with two state-of-the-art MLLMs, Qwen2.5-VL-7B (Bai et al., 2025) and InternVL2.5-7B (Chen et al., 2024), under synthetic shift scenarios. In Table 4, we see that Qwen2.5-VL model shows strong correlations between EMI and RP score, and between EMID and its upper bound with remarkable statistical significance. Although the correlations between EMI and RP score are somewhat weak in the case of InternVL2.5, those are still non-negligible correlations (Schober et al., 2018) with the significance level of 0.05, implying the generality of our framework across different models.
Table 4: Verification of EMI and EMID bound with advanced MLLMs. We conduct correlation analysis between EMI and RP, and between EMID and its upper bound (UB) with Qwen2.5-VL-7B (Bai et al., 2025) and InternVL2.5-7B (Chen et al., 2024), and observe statistically significant correlations.
Model EMI $\Leftrightarrow$ RP EMID $\Leftrightarrow$ UB
Spearman $\rho$ ($p$-val) Kendall $\tau$ ($p$-val) Pearson $r$ ($p$-val)
Qwen2.5-VL 0.767 ($<$0.001) 0.571 ($<$0.001) 0.672 ($<$0.001)
InternVL2.5 0.375 (0.029) 0.273 (0.023) 0.810 ($<$0.001)
EMI analysis on a specialized domain.
Since the proposed information-theoretic measures rely on embeddings from external models, e.g., CLIP and RoBERTa, to easily compute empirical estimates of MI and JSD, one may wonder whether our framework could be applied to a kind of specialized domain datasets such as medical visual question answering. To investigate this, we adopt LLaVA-Med instruction dataset (Li et al., 2024) as a domain-specific open-ended benchmark on the medical images and corresponding questions. We demonstrate, in Table 5, that EMI and EMID upper bound show strong correlation with RP score and EMID, respectively, indicating the broad applicability of our method in a specialized domain as well.
Table 5: Verification of EMI and EMID bound on LLaVA-Med dataset. We conduct correlation analysis between EMI and RP, and between EMID and its upper bound (UB) with LLaVA v1.5 (7B) on open-ended medical domain QA tasks.
EMI $\Leftrightarrow$ RP EMID $\Leftrightarrow$ UB
Spearman $\rho$ ($p$-val) Kendall $\tau$ ($p$-val) Pearson $r$ ($p$-val)
0.718 ($<$0.001) 0.572 ($<$0.001) 0.930 ($<$0.001)
Application: EMID upper bound as a regularization.
Although our framework mainly stands for the performance analysis of MLLMs under distribution shifts at inference time, one can also leverage EMI and/or EMID upper bound at training time. We give a simple example that instantiates the EMID upper bound as an additional regularization term during visual instruction tuning as below,
$\mathbb{E}[H(P_{\theta}(\cdot|\mathbf{x}))]\cdot(D^{\frac{1}{2}}_{\rm JS}(P_{Z_{v}}||\mathcal{N})+D^{\frac{1}{2}}_{\rm JS}(P_{Z_{t}}||\mathcal{N}),$ (11)
where $\mathbf{Z}=(Z_{v},Z_{t})$ denotes an intermediate representation of MLLM given $\mathbf{X}$ and $\mathcal{N}$ is an isotropic Gaussian distribution that has the same dimensionality as $Z_{v}$ and $Z_{t}$. Here, we utilize a non-informative prior distribution $\mathcal{N}$ as an alternative to the target distribution $Q$ (which is usually inaccessible during training) to implicitly penalize representation discrepancy between $P$ and $Q$. The above term encourages intermediate representations of visual and text inputs to be shrunk into a zero-centered Gaussian where the penalizing strength is scaled by an averaged model output entropy across batch samples. Table 6 presents evaluation results of LLaVA v1.5 7B model on LLaVA-Bench COCO (ID) and its synthetic shift variants (V, T, and J Shift) after being instruction tuned on LLaVA-mix-665k subset, where we see our EMID-based regularization effectively improves robustness to shifts while maintaining the ID performance.
Table 6: Visual instruction tuning with EMID upper bound. We use EMID upper bound as a regularization term (Eq. 11) during instruction tuning of LLaVA v1.5 (7B) on a 10% subset of LLaVA-mix-665K w/ and w/o $R$, and report relative preference scores.
Method ID V Shift T Shift J Shift
Baseline 72.7 65.8 68.0 59.6
Baseline w/ $R$ 72.7 66.3 68.3 60.8
@@ 6 Related Works
Recent findings imply that fine-tuning on a relatively small amount of ID datasets hurts the OOD generalization capability of foundation models (Kumar et al., 2022; Wortsman et al., 2022). Although lots of follow-up studies (Goyal et al., 2023; Tian et al., 2023; Oh et al., 2025) including theory-inspired methods (Kumar et al., 2022; Ju et al., 2022; Oh et al., 2024) have been proposed, almost all of them focused on a discriminative model such as CLIP (Radford et al., 2021) for the image classification task. Given the rapidly growing popularity of MLLMs, it is necessary to investigate the reliability of MLLMs under distribution shifts with a tangible formulation. We lay a cornerstone for this.
@@
The information-theoretic view has been steadily adopted to establish evaluation criteria for language model probing (Hewitt et al., 2021), prompt engineering (Sorensen et al., 2022), and rationale evaluation (Chen et al., 2023), alongside learning objectives (Alemi et al., 2016; Chen et al., 2016; Tschannen et al., 2020; Kong et al., 2020; Wang et al., 2021), but relatively unexplored for MLLM yet. We also note some works adopting information-theoretic approaches to analyze models under distribution shifts (Federici et al., 2021; Shui et al., 2022). Although they focused on classification tasks with discriminative models, we established new theorems for MLLM analysis based on a new metric, EMI.